OpenStudy (anonymous):

A hot air balloon ascends 50feet for every 1000 feet travel horizontally.

OpenStudy (anonymous):

What is the problem regarding the balloon and it's enviornment?

OpenStudy (anonymous):

determine the angle of elevation or depression in situation

OpenStudy (anonymous):

Balloon rises 1 foot for every 20 feet of horizontal movement. That makes an angle of \[\frac{180 \text{ArcTan}\left[\frac{1}{20}\right]}{\pi } =2.86241\text{ degrees}\]